If the odds of having H1N1 flu is 0.2, what is the probability of having H1N1 flu? A. 0.2 B. 0.167 C. 4 D. 0.25
Answer #1

Odds = Probability / (1 - Probability)

So,

0.2 = p/(1-p)

2/10 = p/(1-p)

1/5 = p/(1-p)

1 - p = 5p

P = 1/6 = 0.167